Find three consecutive integers such that the sum of three times the first integer, one-third of the second integer, and eight times the third integer is -29.
Answer by stanbon(48558)

Let 1st integer be "x"
2nd is "x+1"
3rd is "x+2"
EQUATION:
3x+(1/3)(x+1)+8(x+2)=-29
Multiply through be "3" to get:
9x+(x+1)+24(x+2)=-87
9x+x+1+24x+48=-87
34x=-136
x=-4
x+1=-3
x+2=-2
Cheers,
Stan H.